# Equivariant Symmetry-Aware Head Pose Estimation for Fetal MRI

### v4

@@ Abstract
We present E(3)-Pose, a novel fast pose estimation method that jointly and explicitly models rotation equivariance and object symmetry. Our work is motivated by the challenging problem of accounting for fetal head motion during a diagnostic MRI scan. We aim to enable automatic adaptive prescription of 2D diagnostic MRI slices with 6-DoF head pose estimation, supported by 3D MRI volumes rapidly acquired before each 2D slice. Existing methods struggle to generalize to clinical volumes, due to pose ambiguities induced by inherent anatomical symmetries, as well as low resolution, noise, and artifacts. In contrast, E(3)-Pose captures anatomical symmetries and rigid pose equivariance by construction, and yields robust estimates of the fetal head pose. Our experiments on publicly available and representative clinical fetal MRI datasets demonstrate the superior robustness and generalization of our method across domains. Crucially, E(3)-Pose achieves state-of-the-art accuracy on clinical MRI volumes, paving the way for clinical translation. Our implementation is available at github.com/ramyamut/E3-Pose.
Project page: ramyamut.github.io/e3-pose/ This work is supported by the the MIT Abdul Latif Jameel Clinic, the MIT-Takeda Program, the Chou Family Transformative Research Fund, NIH NIBIB (R01EB032708, 5T32EB1680), and NIH NICHD (R01HD100009). This work involved human subjects or animals in its research. Approval of all ethical and experimental procedures and protocols was granted by the Institutional Review Board of Boston Children’s Hospital.
@@ 1 Introduction
Figure 1: E(3)-Pose is a rotation-equivariant and symmetry-aware framework for 6-DoF pose estimation. Left: a rapid navigator volume is inserted between every two 2D diagnostic MRI slices. It is used to estimate the fetal head pose to adjust imaging plane prescription in real time. Right: To enable robust performance, the network architecture employs E(3)-equivariant convolutional filters to capture pose equivariance and pseudovectors to account for left-right head symmetry.
@@
Real-time 6-DoF object pose estimation enables a wide range of critical navigation tasks in robotic manipulation [billard2019], autonomous driving [chiu2021], and image-guided surgery [yaniv2006]. Here, we present a novel method for rapid object pose estimation from 3D images based on an architecture that captures pose equivariance and object symmetries by construction. We are motivated by the problem of motion estimation in fetal MRI, where stacks of 2D diagnostic MRI slices are acquired to assess fetal development and detect pathology [glenn2009, gholipour2014] (Appendix A). Due to inter-slice fetal motion, oblique slices (i.e., misaligned with the target anatomical orientation) and spatial coverage gaps in the acquired stack present challenges to radiological assessment and often necessitate re-acquisition of the entire stack (Fig. 1) [malamateniou2013, xu2023, white2010]. One solution is to estimate the fetal head pose from rapid, low-resolution, low signal-to-noise ratio (SNR) 3D MRI “navigator” volumes inserted into the MRI sequence before every 2D slice. The pose estimated from each navigator volume can be used to adaptively adjust the imaging plane of the next 2D slice in real time (Fig. 1) [hess2011, tisdall2012, gagoski2016]. Estimating fetal head pose from navigator volumes is particularly challenging. In addition to low resolution and low SNR, these volumes contain an imaging artifact that presents as a dark plane in the position of the previous (high energy) 2D slice and may obstruct relevant anatomical features (e.g., eyes) that are crucial for pose disambiguation. Moreover, underdeveloped neuroanatomy coupled with low resolution creates a left-right symmetry, introducing pose ambiguities [zhan2013]. Finally, the lack of available navigator data necessitates training on higher resolution, higher SNR, artifact-free research volumes, thus creating cross-domain generalization challenges.
@@
Existing pose estimation methods detect pre-specified anatomical landmarks to compute pose [hoffmann2021, silva2024], but landmark visibility is not guaranteed in clinical data. Another class of methods formulates the problem as rigid registration of a moving volume to a template [moyer2021, billot2024, wang2025]. However, in navigator volumes, artifacts perturb the intensity distributions of moving volumes relative to the template, leading to suboptimal registration. Canonical pose regression with convolutional neural networks (CNNs) can improve robustness in the presence of noise and reduced anatomical visibility [xiang2018, salehi2019, gao2020_2]. Despite their benefits, these methods are prone to overfitting even with data augmentations, and are highly sensitive to pose ambiguities.
@@
We propose E(3)-Pose, a novel framework for accurate and generalizable head pose estimation from fetal brain volumes (Fig. 1). In contrast to standard convolutions that are by construction only equivariant to translations, we leverage an E(3)-equivariant CNN (E(3)-CNN) for pose regression to also account for rotation equivariance [weiler2018, geiger2022]. Moreover, we use pseudovectors [hauser1986] in the learned equivariant rotation parametrization to model left-right head symmetry. E(3)-Pose achieves competitive performance on publicly available research-quality datasets, and state-of-the-art performance on clinically representative datasets. Crucially, our experiments demonstrate that constraining the network architecture to explicitly model physical symmetries provides better generalization to out-of-distribution data and stabilizes pose estimates in ambiguous cases with low anatomical visibility. Overall, while this paper focuses on fetal MRI, our method holds promise for object localization tasks that involve symmetries, limited training data, and real-world applications with noisy sensors.
@@ 2 Related work
Landmark registration involves algorithmic detection of pre-defined anatomical structures to estimate pose relative to a canonical anatomical frame [hill2001, hoffmann2021, silva2024]. To mitigate dependence on visibility of specific anatomy, recent approaches trained a neural network to learn an overcomplete set of landmarks [wang2023, billot2024]. Nevertheless, learned landmarks display sensitivity to perturbations of volume intensities, posing problems when fast acquisitions introduce disruptive artifacts. In contrast, we directly regress pose, circumventing these challenges.
@@
Early pose regression approaches relied on rotation parametrizations that suffer from discontinuities over SO(3), such as axis-angle vectors or quaternions [salehi2019, grassia1998]. It has been shown that high-dimensional, continuous parametrizations offer smoother loss landscapes and robustness to noise [geist2024, zhou2019]. Existing methods that adopt this strategy regress two basis directions [faghihpirayesh2023, zhou2019] or a deformation field [gopinath2024] relative to the object frame. E(3)-Pose similarly regresses a continuous, overcomplete parametrization by using three basis directions of the canonical object frame.
@@
A function is equivariant under a group of transforms if it commutes with every transform in the group. In the context of CNNs, standard convolutions are already equivariant to translations. There has been significant progress toward convolutions that are equivariant under a broader group, including SE(3) and E(3) [weiler2018, geiger2022, cohen2016, cohen2017, weiler2019, thomas2022, worrall2017, cohen2018]. Specifically, these architectures restrict the space of learnable parameters to those that strictly adhere to problem symmetries, rather than learning equivariance through data augmentation. Theoretical analysis shows that equivariant architectures are more sample-efficient, reducing overfitting especially when training data is limited [lawrence2022, sannai2019, bulusu2022]. By leveraging an E(3)-equivariant CNN for pose estimation, E(3)-Pose robustly generalizes to noisy clinical data.
Existing equivariant methods for 6-DoF pose estimation either predict the relative transform between image pairs [moyer2021, wang2022, billot2024, lee2022, zhang2020]; regress a canonical pose via a discontinuous rotation parametrization, such as quaternions [li2021, musallam2022]; or formulate rotation estimation as a classification problem over a discretized set of 3D rotations [howell2023, lee2024, chen2021]. Furthermore, these methods do not explicitly account for symmetric ambiguities in their network architectures. In contrast, E(3)-Pose regresses a continuous output over SO(3) and handles object symmetry in the construction of this parametrization.
Figure 2: Overview of E(3)-Pose. We first train a CNN $\psi$ to segment the object. We estimate translation based on the center-of-mass of the predicted mask and then crop the 3D volumes around this mask. The cropped volumes are fed to an E(3)-CNN $\phi$ trained independently to regress the orthonormal basis of the object frame, parametrized as one pseudovector (red) and two vectors (blue and green). The output is later constrained to represent a rotation matrix by applying SVD and then choosing the pseudovector $\hat{e}_{x}$ direction that results in a proper rotation without reflection (i.e., $\det(M(\hat{R}))=1$).
@@
Rotational and reflectional object symmetries can lead to pose ambiguities. One class of solutions trains a network to implicitly learn object symmetries via probabilistic outputs, and heuristically selects the most probable pose during inference [hodan2020, zhao2023]. Yet, these methods suffer from discontinuities in the network output over the object symmetry group [richterklug2021]. Although training with symmetry-invariant losses partially alleviates this problem [xiang2018], these discontinuities are more robustly eliminated by incorporating object symmetry into the pose parametrization itself. Namely, this strategy regresses an output that remains invariant under the object symmetry group, while maintaining continuity over the quotient group of all non-symmetric input transformations [pitteri2019, saxena2007, richterklug2021]. To handle head pose ambiguities, E(3)-Pose models the basis direction along the axis of left-right symmetry as a pseudovector [hauser1986], which maintains both invariance under left-right reflections and continuity over SO(3). While pseudovectors have been used to model rotation axes in pose estimation [lee2019], our method is the first to use them to address symmetric ambiguities.
@@
Existing implementations use the pose estimated from a single volume acquired at the start of the sequence to prescribe all diagnostic 2D slices in the stack [hoffmann2021, silva2024]. While this approach accurately prescribes slices according to the initial head pose, it does not account for inter-slice head motion. In contrast, E(3)-Pose promises to adjust the imaging plane for every slice in the stack, based on the head pose estimated from the preceding navigator volume.
@@ 3 Preliminaries
Let $f\!:\!\mathbb{R}^{3}\!\rightarrow\!\mathbb{R}^{d}$ be a $d$-dimensional tensor field. Consider 3D Euclidean group actions $g\!=\!g_{t}\circ g_{r}$, which comprise rotations/reflections $g_{r}\!\in\!\text{O(3)}$ and translations $g_{t}\!\in\!\mathbb{R}^{3}$. When the coordinate system is transformed by $g$, the tensor $f(x)$ at point $x\!\in\!\mathbb{R}^{3}$ is modified in two ways: it is moved to the location $g^{-1}x$ and the tensor is independently rotated/reflected by $g_{r}$ [weiler2018, geiger2022]. Formally, we can define an operator $\pi_{f}(\cdot)$ that transforms field $f$ for any $g\!=\!g_{t}\circ g_{r}\!\in\!\text{E(3)}$:
$\pi_{f}(g)f](x)\!=\!\rho_{f}(g_{r})f(g^{-1}x),$ (1)
where $\rho_{f}\!:\!\text{O(3)}\!\rightarrow\!\mathbb{R}^{d\times d}$ is called a representation; it is a matrix acting linearly on the tensor space of $f(x)$ that describes how the tensor independently transforms under $g_{r}$.
@@
A representation $\rho$ and the tensors it transforms are irreducible if there is no change of basis $Q\!\in\!\mathbb{R}^{d\times d}$ such that $Q\rho(g_{r})Q^{-1}$ is block-diagonal for all $g_{r}\!\in\!\text{O(3)}$. Irreducible representations and tensors are characterized by order $l$ and parity, defining how $f(x)$ transforms under proper rotations and inversions (i.e., reflections across all three planes) of the coordinate system, respectively [geiger2022]. Even parity indicates no change in $f(x)$ under inversion, while odd parity indicates a sign change in $f(x)$. While one can define tensors and representations of any order and parity, we focus here on three types: first, scalars $f(x)\!\in\!\mathbb{R}$ (order l=0l\!=\!0, even parity) have representation $\rho^{l=0}_{{\text{even}}}(g_{r})\!=\!1$; second, vectors $f(x)\!\in\!\mathbb{R}^{3}$ (l=1l\!=\!1, odd parity) have representation $\rho^{l=1}_{{\text{odd}}}(g_{r})\!=\!M(g_{r})$, where $M\!:\!\text{O(3)}\!\rightarrow\!\mathbb{R}^{3\times 3}$ is the matrix representation of an orthogonal transformation; and last, pseudovectors $f(x)\!\in\!\mathbb{R}^{3}$ (l=1l\!=\!1, even parity) have representation $\rho^{l=1}_{{\text{even}}}(g_{r})\!=\!(\det M(g_{r}))M(g_{r})$ with a positive determinant to account for invariance under inversions [hauser1986].
@@
E(3)-CNNs can be built using successive equivariant layers, since their composition remains equivariant. Here we describe equivariant convolutional layers (see Appendix B for pooling and non-linearities). For enhanced computation and time efficiency, state-of-the-art implementations of such CNNs operate on irreducible fields [weiler2018, geiger2022]. Let $\kappa_{l_{f}p_{f},l_{h}p_{h}}(x)\in\mathbb{R}^{(2l_{h}+1)\times(2l_{f}+1))}$ be a 3D convolution kernel transforming field $f$ (order $l_{f}$, parity $p_{f}$) to field $h$ (order $l_{h}$, parity $p_{h}$). Then, $\kappa$ must satisfy Eq. (2) to maintain equivariance, resulting in the solution
@@
where vec indicates vectorization, $w_{jm}$ are coefficients, $\{\varphi_{m}\}$ are radially symmetric basis functions, $Q_{j}\!\in\!\mathbb{R}^{(2l_{h}+1)(2l_{f}+1)\times(2j+1)}$ is a change of basis, $Y_{j}\!:\!S^{2}\rightarrow\mathbb{R}^{2j+1}$ are the degree-$j$ spherical harmonics, and $J\!=\!\{j\!:\!|l_{f}-l_{h}|\!\leq\!j\!\leq\!|l_{f}+l_{h}|,j\text{ is even}\}$ if $p_{f}\!=\!p_{h}$ or $J\!=\!\{j\!:\!|l_{f}-l_{h}|\!\leq\!j\!\leq\!|l_{f}+l_{h}|,j\text{ is odd}\}$ if $p_{f}\!\neq\!p_{h}$ [weiler2018, geiger2022]. E(3)-CNNs are thus trained by learning the coefficients $w_{jm}$ rather than the kernel values themselves.

### v5

@@ Abstract
We present E(3)-Pose, a novel fast pose estimation method that jointly and explicitly models rotation equivariance and object symmetry. Our work is motivated by the challenging problem of accounting for fetal head motion during a diagnostic MRI scan. We aim to enable automatic adaptive prescription of 2D diagnostic MRI slices with 6-DoF head pose estimation, supported by 3D MRI volumes rapidly acquired before each 2D slice. Existing methods struggle to generalize to clinical volumes, due to pose ambiguities induced by inherent anatomical symmetries, as well as low resolution, noise, and artifacts. In contrast, E(3)-Pose captures anatomical symmetries and rigid pose equivariance by construction, and yields robust estimates of the fetal head pose. Our experiments on publicly available and representative clinical fetal MRI datasets demonstrate the superior robustness and generalization of our method across domains. Crucially, E(3)-Pose achieves state-of-the-art accuracy on clinical MRI volumes, supporting future clinical translation. Our implementation is publicly available [^1] Code: https://github.com/MedicalVisionGroup/E3-Pose.
Project page: ramyamut.github.io/e3-poseThis work is supported by the the MIT Abdul Latif Jameel Clinic, the MIT-Takeda Program, the Chou Family Transformative Research Fund, NIH NIBIB (R01EB032708, 5T32EB1680), and NIH NICHD (R01HD100009). This work involved human subjects or animals in its research. Approval of all ethical and experimental procedures and protocols was granted by the Institutional Review Board of Boston Children’s Hospital.
@@ 1 Introduction
Figure 1: E(3)-Pose is a rotation-equivariant and symmetry-aware framework for 6-DoF pose estimation. Left: a rapid navigator volume is inserted between every two 2D diagnostic MRI slices. Our aim is to estimate the fetal head pose to adjust imaging plane prescription in real time. Right: To enable robust performance, the network architecture employs E(3)-equivariant convolutional filters to capture pose equivariance and pseudovectors to account for left-right head symmetry.
@@
Real-time 6-DoF object pose estimation enables a wide range of critical navigation tasks in robotic manipulation [2], autonomous driving [9], and image-guided surgery [78]. Here, we present a novel method for rapid object pose estimation from 3D images based on an architecture that captures pose equivariance and object symmetries by construction. We are motivated by the problem of motion estimation in fetal MRI, where stacks of 2D diagnostic MRI slices are acquired to assess fetal development and detect pathology [24, 23] (Appendix A). Due to inter-slice fetal motion, oblique slices and spatial coverage gaps in the acquired stack present challenges to radiological assessment and often necessitate re-acquisition of the entire stack [48, 76, 72]. One solution is to estimate the fetal head pose from rapid, low-resolution, low signal-to-noise ratio (SNR) 3D MRI “navigator” volumes inserted into the MRI sequence before every 2D slice. The aim is to use the pose estimated from each navigator volume to adaptively adjust the imaging plane of the next slice (Fig. 1) [30, 66, 18]. Estimating fetal head pose from navigator volumes is particularly challenging. In addition to low resolution and SNR, these volumes contain an imaging artifact that presents as a dark plane in the position of the previous (high energy) slice and may obstruct relevant anatomical features (e.g., eyes) that are crucial for pose disambiguation. Moreover, underdeveloped neuroanatomy coupled with low resolution creates an approximate left-right symmetry, introducing pose ambiguities [79]. Finally, the lack of available navigator data necessitates training on higher resolution, higher SNR, artifact-free research volumes, thus creating cross-domain generalization challenges.
@@
Several existing methods detect pre-specified anatomical landmarks to compute pose [33, 64], but landmark visibility is not guaranteed in clinical data. Another class of methods formulates the problem as rigid registration of a moving volume to a template [52, 3, 68]. However, in navigator volumes, artifacts perturb the intensity distributions of moving volumes relative to the template, leading to suboptimal registration. Canonical pose regression with convolutional neural networks (CNNs) can improve robustness to reduced anatomical visibility and imaging artifacts [75, 60, 20]. Despite their benefits, these methods are prone to overfitting even with data augmentations, and are highly sensitive to pose ambiguities.
@@
We propose E(3)-Pose, a novel framework for accurate and generalizable head pose estimation from fetal brain volumes (Fig. 1). In contrast to standard convolutions that are by construction only equivariant to translations, we leverage an E(3)-equivariant CNN (E(3)-CNN) for pose regression to also account for rotation equivariance [71, 21]. Moreover, we use pseudovectors [29] in the learned equivariant rotation parametrization to model left-right head symmetry. E(3)-Pose achieves competitive performance on publicly available research-quality datasets, and state-of-the-art performance on clinically representative datasets. Crucially, our experiments demonstrate that constraining the network architecture to explicitly model physical symmetries provides better generalization to out-of-distribution data and stabilizes pose estimates in ambiguous cases with low anatomical visibility. Overall, while this paper focuses on fetal MRI, our method holds promise for object localization tasks that involve symmetries, limited training data, and real-world applications with noisy sensors.
@@ 2 Related work
Landmark registration involves algorithmic detection of pre-defined anatomical structures to estimate pose relative to a canonical anatomical frame [31, 33, 64]. To mitigate dependence on visibility of specific anatomy, recent approaches trained a neural network to learn an overcomplete set of landmarks [67, 3]. Nevertheless, learned landmarks display sensitivity to perturbations of volume intensities, posing problems when fast acquisitions introduce disruptive artifacts. In contrast, we directly regress pose, circumventing these challenges.
@@
Early pose regression approaches relied on rotation parametrizations that suffer from discontinuities over SO(3), such as axis-angle vectors [60, 27]. It has been shown that high-dimensional, continuous parametrizations offer smoother loss landscapes and robustness to noise [22, 82]. Existing methods that adopt this strategy regress two basis directions [16, 82] or a deformation field [25] of the canonical object frame relative to the input volume. E(3)-Pose similarly regresses a continuous, overcomplete parametrization by using three basis directions.
@@
A function is equivariant under a group of transforms if it commutes with every transform in the group. In the context of CNNs, standard convolutions are already equivariant under translations. There has been significant progress toward convolutions that are equivariant under a broader group, including SE(3) and E(3) [71, 21, 12, 13, 70, 65, 74, 11]. Specifically, these architectures restrict the space of learnable parameters to those that strictly adhere to inherent symmetries, rather than learning equivariance through data augmentation. Theoretical analysis shows that equivariant architectures are more sample-efficient, reducing overfitting especially when training data is limited [41, 61, 6]. By leveraging an E(3)-equivariant CNN for pose estimation, E(3)-Pose robustly generalizes to noisy clinical data.
Existing equivariant methods for 6-DoF pose estimation either predict the relative transform between image pairs [52, 69, 3, 44, 80] or formulate rotation estimation as a classification problem over a discretized set of 3D rotations [46, 53, 35, 43, 8]. Furthermore, these methods do not explicitly account for symmetric ambiguities in their network architectures. In contrast, E(3)-Pose regresses a continuous output over SO(3) and handles object symmetry in the construction of this parametrization.
Figure 2: Overview of E(3)-Pose. We first train a CNN $\psi$ to segment the object. We estimate translation based on the center-of-mass of the predicted mask and then crop the 3D volumes around this mask using a 40% margin. The cropped volumes are fed to an E(3)-CNN $\phi$ trained independently to regress the orthonormal basis of the object frame, parametrized as one pseudovector (red) and two vectors (blue and green). The output is later constrained to represent a rotation matrix by applying SVD and then choosing the pseudovector $\hat{e}_{x}$ direction that results in a proper rotation without reflection (i.e., $\det(M(\hat{R}))=1$).
@@
Rotational and reflectional object symmetries can lead to pose ambiguities. One class of solutions trains a network to implicitly learn object symmetries via probabilistic outputs, and heuristically selects the most probable pose during inference [32, 81]. Yet, these methods suffer from discontinuities in the network output over the object symmetry group [58]. Although training with symmetry-invariant losses partially alleviates this problem [75], these discontinuities are more robustly eliminated by incorporating object symmetry into the pose parametrization itself. Namely, this strategy regresses an output that remains invariant under the object symmetry group, while maintaining continuity over the group of all non-symmetric input transformations [56, 62, 58]. To handle head pose ambiguities, E(3)-Pose models the basis direction along the axis of left-right symmetry as a pseudovector [29], which maintains both invariance under left-right reflections and continuity over SO(3). While pseudovectors have been used to model rotation axes in pose estimation [42], our method is the first to use them to address symmetric ambiguities.
@@
Existing implementations use the pose estimated from a single volume acquired at the start of the sequence to prescribe all slices in the stack [33, 64]. While this approach accurately prescribes slices according to the initial head pose, it does not account for inter-slice motion. In contrast, E(3)-Pose aims to adjust the imaging plane for every slice in the stack, based on the head pose estimated from the preceding navigator volume.
@@ 3 Preliminaries
Let $f\!:\!\mathbb{R}^{3}\!\rightarrow\!\mathbb{R}^{d}$ be a $d$-dimensional tensor field. Consider 3D Euclidean group actions $g\!=\!g_{t}\circ g_{r}$, which comprise rotations/reflections $g_{r}\!\in\!\text{O(3)}$ and translations $g_{t}\!\in\!\mathbb{R}^{3}$. When the coordinate system is transformed by $g$, the tensor $f(x)$ at point $x\!\in\!\mathbb{R}^{3}$ is modified in two ways: it is moved to the location $g^{-1}x$ and the tensor is independently rotated/reflected by $g_{r}$ [71, 21]. Formally, we can define an operator $\pi_{f}(g)f](x)\!=\!\rho_{f}(g_{r})f(g^{-1}x)$ that transforms field $f$ for any $g\!=\!g_{t}\circ g_{r}\!\in\!\text{E(3)}$. In this formulation, $\rho_{f}\!:\!\text{O(3)}\!\rightarrow\!\mathbb{R}^{d\times d}$ is called a representation; it is a matrix acting linearly on the tensor space of $f(x)$ that describes how the tensor independently transforms under $g_{r}$.
@@
A representation $\rho$ and the tensors it transforms are irreducible if there is no change of basis $Q\!\in\!\mathbb{R}^{d\times d}$ such that $Q\rho(g_{r})Q^{-1}$ is block-diagonal for all $g_{r}\!\in\!\text{O(3)}$. Irreducible representations and tensors are characterized by order $l$ and parity, defining how $f(x)$ transforms under proper rotations and inversion (i.e., reflection across all three planes) of the coordinate system, respectively [21]. Even parity indicates no change in $f(x)$ under inversion, while odd parity indicates a sign change in $f(x)$. While one can define tensors and representations of any order and parity, we focus here on three types: first, scalars $f(x)\!\in\!\mathbb{R}$ (order l=0l\!=\!0, even parity) have representation $\rho^{l=0}_{{\text{even}}}(g_{r})\!=\!1$; second, vectors $f(x)\!\in\!\mathbb{R}^{3}$ (l=1l\!=\!1, odd parity) have representation $\rho^{l=1}_{{\text{odd}}}(g_{r})\!=\!M(g_{r})$, where $M\!:\!\text{O(3)}\!\rightarrow\!\mathbb{R}^{3\times 3}$ is the matrix representation of an orthogonal transformation; and last, pseudovectors $f(x)\!\in\!\mathbb{R}^{3}$ (l=1l\!=\!1, even parity) have representation $\rho^{l=1}_{{\text{even}}}(g_{r})\!=\!(\det M(g_{r}))M(g_{r})$ with a positive determinant to account for invariance under inversion [29].
@@
E(3)-CNNs can be built using successive equivariant layers, since their composition remains equivariant. Here we describe equivariant convolutional layers (see Appendix B for pooling and non-linearities). For enhanced computation and time efficiency, state-of-the-art implementations of such CNNs operate on irreducible fields [71, 21]. Let $\kappa_{l_{f}p_{f},l_{h}p_{h}}(x)\in\mathbb{R}^{(2l_{h}+1)\times(2l_{f}+1))}$ be a 3D convolution kernel transforming field $f$ (order $l_{f}$, parity $p_{f}$) to field $h$ (order $l_{h}$, parity $p_{h}$). Then, $\kappa$ must satisfy Eq. (1) to maintain equivariance, resulting in the solution
@@
where vec indicates vectorization, $w_{jm}$ are coefficients, $\{\varphi_{m}\}$ are radially symmetric basis functions, $Q_{j}\!\in\!\mathbb{R}^{(2l_{h}+1)(2l_{f}+1)\times(2j+1)}$ are changes of basis, $Y_{j}\!:\!S^{2}\rightarrow\mathbb{R}^{2j+1}$ are the degree-$j$ spherical harmonics, and $J\!=\!\{j\!:\!|l_{f}-l_{h}|\!\leq\!j\!\leq\!|l_{f}+l_{h}|,j\text{ is even}\}$ if $p_{f}\!=\!p_{h}$ or $J\!=\!\{j\!:\!|l_{f}-l_{h}|\!\leq\!j\!\leq\!|l_{f}+l_{h}|,j\text{ is odd}\}$ if $p_{f}\!\neq\!p_{h}$ [71, 21]. E(3)-CNNs are thus trained by learning the coefficients $w_{jm}$ rather than the kernel values themselves.